# Some phenomenological aspects of a quantum-corrected Reissner–Nordström black hole: quasi-periodic oscillations, scalar perturbations and thermal fluctuations

### v1

@@ VII Thermal fluctuations
We now investigate the effect of thermal fluctuations on the BH entropy. Within the Euclidean quantum gravity framework, the partition function is obtained by performing a Wick rotation of the time coordinate and expanding the gravitational action about the saddle-point (classical) geometry [57]. In this approach, thermal (or quantum) fluctuations around equilibrium lead to corrections to the leading semiclassical entropy. In the saddle-point approximation, the leading contribution reproduces the Bekenstein–Hawking entropy, $S=A/4=\pi r_{h}^{2}$, while Gaussian fluctuations around equilibrium generate subleading corrections. For a canonical ensemble, incorporating thermal fluctuations leads to a corrected entropy of the form [69, 58]:
@@
which is analogue to that for RN BH case.
@@ Acknowledgements.
Data Availability Statement
This manuscript has no associated data.
Conflict of Interest
Authors declare(s) no conflict of interest.

### v2

@@ VII Thermal fluctuations
We now investigate the effect of thermal fluctuations on the BH entropy. Within the Euclidean quantum gravity framework, the partition function is obtained by performing a Wick rotation of the time coordinate and expanding the gravitational action about the saddle-point (classical) geometry [63]. In this approach, thermal (or quantum) fluctuations around equilibrium lead to corrections to the leading semiclassical entropy. In the saddle-point approximation, the leading contribution reproduces the Bekenstein-Hawking entropy, $S=A/4=\pi r_{h}^{2}$, while Gaussian fluctuations around equilibrium generate subleading corrections. For a canonical ensemble, incorporating thermal fluctuations leads to a corrected entropy of the form [74, 64]:
@@
which is similar to that for the RN black hole.
Figure 15: Behavior of the corrected entropy expression as a function of horizon $r_{h}$. Here $Q/M=0.2$.